need to know how to graph the equation for the following problem
1.y=2x-5
2.y=2/3x+-4
3.4x-5y=-20
4.Find the slope of the line passing through (-7,3) and (2,6)
does the line rise or fall from left to right
5.one line passes through(7,20 and (4,4). Another line passes through (0,10) and (-2,7).without graphing determine whether the lines are parallel, perpendicular or neither, I would like to also see this in a graph please
Thank You

Answer by Alan3354(69443)

need to know how to graph the equation for the following problem
1.y=2x-5
2.y=2/3x+-4
3.4x-5y=-20
----
These are straight lines, so you need 2 points.
Pick 2 values for x and find y.
Plot the points.
Draw a line thru them.
=======================
4.Find the slope of the line passing through (-7,3) and (2,6)
does the line rise or fall from left to right
--
Slope = diffy/diffx
=============================
5.one line passes through(7,20 and (4,4). Another line passes through (0,10) and (-2,7).without graphing determine whether the lines are parallel, perpendicular or neither, I would like to also see this in a graph please
Find the 2 slopes, m1 & m2.
If they're equal --> parallel
If m1 = -1/m1 they're perpendicular.
o/w neither.
-------------
DL the FREE graph software at
www.padowan.dk